Equivalent Units of Conversion Costs

The Filling Department of Eve Cosmetics Company had 5,500 ounces in beginning work in process inventory (70% complete). During the period, 45,700 ounces were completed. The ending work in process inventory was 2,700 ounces (80% complete).

What are the total equivalent units for conversion costs?

If required, round to the nearest unit.
Answer #1

Equivalent unit of conversion Under Weighted average method = Unit transferred out+(Ending work in process*Percent completion)

                                                = 45700+(2700*80%)

Equivalent unit of conversion under weighted average method = 47860 ounces

Note : Question does not specify method so i have with both method

Equivalent unit under FIFO method = (5500*30%+40200*100%+2700*80%) = 44010 Ounces